# Interpret Your Decision: Logical Reasoning Regularization for Generalization in Visual Classification

### v1

@@ 1 Introduction
Intuitively, the complexity reduction achieved by L-Reg stems from its ability to filter out redundant features or semantics, focusing instead on the minimal yet sufficient semantics for classification - defined as semantic support in Definition 3.2, where the interpretability also emerges. This filtering feature benefits the generalization when there is a domain shift in data where the domain-dependent features are ignored for classification. Moreover, it further promotes generalization when unlabeled data from the unknown classes is present. If such data lacks the semantic support associated with known classes, it is then classified as belonging to an unknown class. These capabilities equip L-Reg with explicit interpretability. As Fig. 1 shows, with L-Reg, the model can identify the unknown class ‘person’, and pinpoint faces which are the crucial features for classifying this category. In contrast, the model trained solely with $L_{2}$ (without L-Reg) focuses on the ambiguous features for classification.
Rigorous theoretical analysis and experimental results validate that L-Reg yields better generalization across diverse scenarios. Specifically, L-Reg facilitates better performance under the aforementioned multi-domain generalization and generalized category discovery tasks, whose settings are presented in Fig. 2 (a)(b). Furthermore, to evaluate L-Reg’s robustness, we introduce a more complex real-world scenario, as shown in Fig. 2 (c), where unlabeled images may not only belong to unknown classes, but also originate from unseen domains. Even in this challenging context, L-Reg is still able to consistently demonstrate notable improvements in generalization, underscoring its practical utility and effectiveness.
@@ 2 Preliminaries and generalization settings for visual classification
Let the target model $f^{*}:f^{*}(X,Y):X\to Y$, can generalize across both seen and unseen sets $X,Y$. Denote its trainable ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f}$, which is only trained on the seen sets. The generalization loss for the unseen sets is defined as:
$GL({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f},f^{*},(X_{u},Y_{u}))=\mathbb{E}_{(x,y)\in({X}_{u},{Y}_{u})}||f(x,y)-f^{*}(x,y)||_{2}.$ (1)
Figure 3: Visualizations of classifiers’ weights form models trained using GMDG on PACS dataset without and with L-Reg under mDD+GCD setting, respectively. Both experiments share the same hyper-parameters using Regnety-16g backbone, except the latter uses additional L-Reg.
@@ 3 Logical regularization for generalization in image classification
$L_{L-Reg}\!\!=\!\!\frac{1}{M}\sum_{i=1}^{M}\left[\sum_{j=1}^{K}[\sigma(Z_{i}^{T}\hat{Y}_{j})\log\sigma(Z_{{i}}^{T}\hat{Y}_{j})]-[\frac{1}{K}\sum_{j=1}^{K}\sigma(Z^{T}_{i}\hat{Y}_{j}))\log(\frac{1}{K}\sum_{j=1}^{K}\sigma(Z^{T}_{i}\hat{Y}_{j})]\right]\!,$ (3)
@@ 3.1 Logical framework for visual classification
For clarity, we specify $\mathcal{L}_{(X_{s},Y_{s})}=\left\langle F_{(X_{s},Y_{s})},D,\models_{(X_{s},Y_{s})},h,\vdash_{(h(X),Y)}\right\rangle$ as the logic formed on the given $X,Y$ sets. With the goal for logic to generalize across a broader scenario and provide extrapolation across all possible formulas in $\mathcal{L}$, a ‘good general’ logic $\mathcal{L}^{*}$ should be derived from $\mathcal{L}$ through the feature extractor $g$:
@@
Importantly, as a ‘good general’ logic, $F_{(g(X_{s}),Y_{s})}$ in $\mathcal{L}^{*}$ should be the atomic formulas, i.e., the tuple of terms with a predicate, which makes that $\vdash_{(h\circ g(X_{u}),Y_{u})}=\models_{(g(X_{s}),Y_{s})}$ still holds. Additionally, $\vdash_{(h\circ g(X),Y)}=\models_{(g(X_{s}),Y_{s})}$ in Eq. 6 can be safely omitted in the rest of the paper. Please see more details about the conditions of the good general logic in Appendix B.
@@ 3.2 Constructing atomic formulas using L-Reg
Interpretability of semantic supports roots in forming atomic formulas. The atomic formula $\mathcal{A}^{y}$ is of the form $h(f(g(x)),y,d)$ or $h(f(z,y),d)$. Our aim is to find the good (most) general $\mathcal{A}^{y*}\in\mathcal{A}^{y}$ for $y$ class from which the interpretability of L-Reg is derived. Consider $\mathcal{A}^{y}_{1},\mathcal{A}^{y}_{2}\in\mathcal{A}^{y}$, if $\mathcal{A}^{y}_{1}$ is more general than $\mathcal{A}^{y}_{2}$, there will be a substitution $\psi$ such that $\mathcal{A}^{y}_{1}\psi=\mathcal{A}^{y}_{2}$ [52]. $\mathcal{A}^{y*}$ should meet $\mathcal{A}^{y*}\psi=\mathcal{A}^{y}_{i}\in\mathcal{A}^{y}$, which infers that $\gamma^{y}\psi=z^{y}$ (cf. Eq. 9) for predication of $y$ where $\gamma^{y}$ is the semantic support. Note here that the form of $\mathcal{A}^{y}$ is constructed for $y\in Y$, i.e., predicate whether the sample belongs to the $y$ class. Considering multiple classes $y_{i},y_{j}\in Y,i\neq j$, it has $\mathcal{A}^{y_{i}*}\neq\mathcal{A}^{y_{j}*}$ thus $\gamma^{y_{i}}\neq\gamma^{y_{j}}$ (cf. Paper Eq.10), which constrains that different minimal semantic supports should be used for predicting different classes. The interpretability of L-Reg is based on $\mathcal{A}^{y*}$, compelling the model to use distinct minimal semantic supports for each class. These minimal semantic supports can be interpreted as the most critical features for efficient prediction. For example, as shown in Paper Fig.1, the model with L-Reg has learned the facial features of the person class (see more examples in Appendix Figs. 7, 8, 9, 10, 11, 12, 8 and 12), forming the (informal) atomic formula $h(f(\text{has a human face},\text{is person}),d\in D)\rightarrow\text{True}$.
@@ 4 L-Reg under different generalization settings
Test domain PACS VLCS OfficeHome TerraInc DomainNet Avg.
@@ 5.3 Experiments on mDG + GCD
L-Reg forms atomic formulas and improves interpretability. Furthermore, Fig. 5 provides visual insights into the behavior of models trained with L-Reg. Evidently, these models tend to focus on minimal semantics sufficient for class distinctions. For the known classes, the efficacy of L-Reg can be intuitively understood as extracting the minimal semantic supports for a given class label. As examples shown here, For instance, the presence of a guitar’s fingerboard, even in unseen domains, helps classify a sample as belonging to the guitar category, whose informal form can be denoted as $h(f(\text{has fingerboard},\text{is guitar}),d\in D)\to\text{True}$. For all known classes, samples with these minimal semantic supports are recognized accordingly. In contrast, if a sample lacks these minimal supports for any known class, it is very likely categorized as an unknown class. This behavior stems from Paper Eq.10 which ensures $\mathcal{A}^{y_{i}*}\neq\mathcal{A}^{y_{j}*}$ through constraining $\gamma^{y_{i}}\neq\gamma^{y_{j}}$. L-Reg further enhances the model’s ability to identify minimal supports for unknown classes by filtering out co-covariant features associated with other classes and thus generalizing to unseen domains. Therefore, the very interpretable features for unknown classes from unseen domains can be extracted using L-Reg. Fig. 5 (right side) demonstrates that the model with L-Reg can even extract facial features for the unknown person class and can generalize this to the unseen domain. Similarly, here we obtain an (informal) atomic formula as $h(f(\text{has a face},\text{is person}),d\in D)\to\text{True}$.
@@ Appendix B Details of the logical framework for visual classification task
    Any bijection $f:P\rightarrow Q$ that extends to a bijection between the tautological formula algebras of $\mathcal{L}^{P}$ and $\mathcal{L}^{Q}$ induces an isomorphism between $\mathcal{L}^{P}$ and $\mathcal{L}^{Q}$.
@@ Appendix E Apply L-Reg to ERM Baseline for mDG
To further validate L-Reg’s efficacy for mDG, we use ERM as the baseline on the TerraInc dataset. For a fair comparison, all experiments share the same hyperparameter settings and use the Regnety-16gf backbone. Original ERM results are also included alongside our reproduced results. The results in Table 8 reveal that ERM with L-Reg significantly improves mDG performance (from 49.9% to 52.9%).
@@ Appendix F Compare L-Reg with more regularization terms
We also compare L-Reg with other regularization terms: The Ortho-Reg - the orthogonality regularization that constrains the independence of each dimension of the semantic feature $z$; and Sparsity - implemented as Bernoulli Sample of the latent features from the sparse linear concept discovery models [41] on our used PIM backbone. Table 6&Table 7 demonstrate that L-Reg outperforms Ortho-Reg and Sparsity.
@@ Appendix G Limitation of L-Reg and possible solutions
As analyzed and discussed in the paper, L-Reg is based on the precondition that each dimension of the latent features represents an independent semantic. Thus, improper semantic features which do not meet this precondition may lead to sub-optimal results. To validate this hypothesis, we test L-Reg by reinforcing independence with Ortho-Reg. MDG results in Table 8 and GCD results in Table 6&Table 7 show that combining L-Reg with Ortho-Reg leads to further improvements, whereas Ortho-Reg alone may not guarantee improvements. This suggests a direction for future work.
@@
Table 8: Results of mDG: Results of using ERM as the baseline. We use the ERM method as the baseline to test L-Reg’s efficacy. Ortho-Reg: orthogonality regularization. This table includes results: (1) The improved performance of L-Reg on ERM baseline. (2) Comparison between L-Reg and the Ortho-Reg on ERM baseline. (3) Using L-Reg and Ortho-Reg together yields further promotion, validating our ‘improper $z$’ hypothesis in the Paper limitation part. The used dataset is TerraInc. All experiments share the same hyperparameters except the added regularization term. Each group of experiments is run with seeds [0,1,2], and the averaged results for each domain and additionally with the standard deviation (Std) are reported.
@@ H.1 Multi-domain generalization
Datasets. We use PACS (4 domains, 9,991 samples, $7$ classes) [32], VLCS ($4$ domains, $10,729$ samples, $5$ classes) [18], OfficeHome (4 domains, 15,588 samples, 65 classes) [55], TerraIncognita (TerraInc, $4$ domains, $24,778$ samples, $10$ classes) [7], and DomainNet (6 domains, 586,575 samples, 345 classes) [42].
@@
TerraInc Location 100 Location 38 Location 43 Location 46 Avg.
@@ H.3 Combination of multi-domain generalization and generalized category discovery
PACS HomeOffice VLCS TerraInc DomainNet
@@
Table 23: MDG+GCD results: accuracy scores of each domain in TerraInc dataset.
TerraInc Avg art_painting cartoon photo sketch

### v3

@@ 1 Introduction
Intuitively, the complexity reduction achieved by L-Reg stems from its ability to filter out redundant features or semantics, focusing instead on the minimal yet sufficient semantics for classification - defined as semantic support in Definition 3.2, where the interpretability also emerges. This filtering feature benefits the generalization when there is a domain shift in data where the domain-dependent features are ignored for classification. Moreover, it further promotes generalization when unlabeled data from the unknown classes is present. If such data lacks the semantic support associated with known classes, it is then classified as belonging to an unknown class, and its corresponding semantic supports are extracted. These capabilities equip L-Reg with explicit interpretability. As Fig. 1 shows, with L-Reg, the model can identify the unknown class ‘person’, and pinpoint faces which are the crucial features for classifying this category. In contrast, the model trained solely with $L_{2}$ (without L-Reg) focuses on the ambiguous features for classification.
Rigorous theoretical analysis and experimental results validate that L-Reg yields better generalization across diverse scenarios. Specifically, L-Reg facilitates better performance under the aforementioned multi-domain generalization and generalized category discovery tasks, whose settings are presented in Fig. 2 (a)(b). Furthermore, to evaluate L-Reg’s robustness, we introduce a more complex real-world scenario, as shown in Fig. 2 (c), where unlabeled images may not only belong to unknown classes, but also originate from unseen domains. Even in this challenging context, L-Reg is still able to consistently demonstrate notable improvements in generalization, underscoring its practical utility and effectiveness. Our code is available at https://github.com/zhaorui-tan/L-Reg_NeurIPS24.
@@ 2 Preliminaries and generalization settings for visual classification
Let the target model $f^{*}:f^{*}(X,Y):X\to Y$, can generalize across both seen and unseen sets $X,Y$. Denote its trainable ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}f}$, which is only trained on the seen sets. The generalization loss for the unseen sets is defined as:
$GL({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}f},f^{*},(X_{u},Y_{u}))=\mathbb{E}_{(x,y)\in({X}_{u},{Y}_{u})}||f(x,y)-f^{*}(x,y)||_{2}.$ (1)
Figure 3: Visualizations of classifiers’ weights form models trained using GMDG on PACS dataset without and with L-Reg under mDG+GCD setting, respectively. Both experiments share the same hyper-parameters using Regnety-16g backbone, except the latter uses additional L-Reg.
@@ 3 Logical regularization for generalization in image classification
$L_{L-Reg}\!\!=\!\!\frac{1}{M}\sum_{i=1}^{M}\left[\sum_{j=1}^{K}[\sigma(Z_{i}^{T}\hat{Y}_{j})\log\sigma(Z_{{i}}^{T}\hat{Y}_{j})]-[\frac{1}{K}\sum_{j=1}^{K}\sigma(Z^{T}_{i}\hat{Y}_{j})\log(\frac{1}{K}\sum_{j=1}^{K}\sigma(Z^{T}_{i}\hat{Y}_{j})]\right]\!,$ (3)
@@ 3.1 Logical framework for visual classification
For clarity, we specify $\mathcal{L}_{(X_{s},Y_{s})}=\left\langle F_{(X_{s},Y_{s})},D,\models_{(X_{s},Y_{s})},h,\vdash_{(h(X),Y)}\right\rangle$ as the logic formed on the given $X,Y$ sets. With the goal for logic to generalize across a broader scenario and provide extrapolation across all possible formulas in $\mathcal{L}$, a good general logic $\mathcal{L}^{*}$ should be derived from $\mathcal{L}$ through the feature extractor $g$:
@@
Importantly, as a good general logic, $F_{(g(X_{s}),Y_{s})}$ and $h$ in $\mathcal{L}^{*}$ should form the atomic formulas, i.e., the tuple of terms with a predicate: $h\circ g(x)\text{ belongs/not belongs to class }y\text{ in domain }d\to Ture/False,\;\text{ where }x,y,d\in X,Y,D$, which makes that $\vdash_{(h\circ g(X_{u}),Y_{u})}=\models_{(g(X_{s}),Y_{s})}$ still holds. We simply denote one atomic formula in the form of $h(g(x),y,d)$ mapping to binary values. Additionally, $\vdash_{(h\circ g(X),Y)}=\models_{(g(X_{s}),Y_{s})}$ in Eq. 6 can be safely omitted in the rest of the paper. Please see more details about the conditions of the good general logic in Appendix B.
@@ 3.2 Constructing atomic formulas using L-Reg
Interpretability of semantic supports roots in forming atomic formulas. The atomic formula $\mathcal{A}^{y}$ is of the form $h(g(x),y,d)$. Our aim is to find the good (most) general $\mathcal{A}^{y*}\in\mathcal{A}^{y}$ for $y$ class from which the interpretability of L-Reg is derived. Consider $\mathcal{A}^{y}_{1},\mathcal{A}^{y}_{2}\in\mathcal{A}^{y}$, if $\mathcal{A}^{y}_{1}$ is more general than $\mathcal{A}^{y}_{2}$, there will be a substitution $\psi$ such that $\mathcal{A}^{y}_{1}\psi=\mathcal{A}^{y}_{2}$ [52]. $\mathcal{A}^{y*}$ should meet $\mathcal{A}^{y*}\psi=\mathcal{A}^{y}_{i}\in\mathcal{A}^{y}$, which infers that $\gamma^{y}\psi=z^{y}$ (cf. Eq. 9) for predication of $y$ where $\gamma^{y}$ is the semantic support. Note here that the form of $\mathcal{A}^{y}$ is constructed for $y\in Y$, i.e., predicate whether the sample belongs to the $y$ class. Considering multiple classes $y_{i},y_{j}\in Y,i\neq j$, it has $\mathcal{A}^{y_{i}*}\neq\mathcal{A}^{y_{j}*}$ thus $\gamma^{y_{i}}\neq\gamma^{y_{j}}$ (cf. Eq. 10), which constrains that different minimal semantic supports should be used for predicting different classes. The interpretability of L-Reg is based on $\mathcal{A}^{y*}$, compelling the model to use distinct minimal semantic supports for each class. These minimal semantic supports can be interpreted as the most critical features for efficient prediction. For example, as shown in Fig. 1, the model with L-Reg has learned the facial features of the person class (see more examples in Appendix Figs. 7, 8, 9, 10, 11, 12, 8 and 12), forming the (informal) atomic formula h(has a $h(\text{has a human face},\text{is person},d\in D)\rightarrow\text{True}$. Similarly, it also leads to h(not has a $h(\text{not has a human face},\text{is person},d\in D)\rightarrow\text{False}$.
@@ 4 L-Reg under different generalization settings
Test domain PACS VLCS OfficeHome TerraIncognita DomainNet Avg.
@@ 5.3 Experiments on mDG + GCD
L-Reg forms atomic formulas and improves interpretability. Furthermore, Fig. 5 provides visual insights into the behavior of models trained with L-Reg. Evidently, these models tend to focus on minimal semantics sufficient for class distinctions. For the known classes, the efficacy of L-Reg can be intuitively understood as extracting the minimal semantic supports for a given class label. For instance, the presence of a guitar’s fingerboard, even in unseen domains, helps classify a sample as belonging to the guitar category, whose informal forms can be denoted as $h(\text{has fingerboard},\text{is guitar},d\in D)\to\text{True}$ and h(not $h(\text{not has fingerboard},\text{is guitar},d\in D)\to\text{False}$. For all known classes, samples with these minimal semantic supports are recognized accordingly. In contrast, if a sample lacks these minimal supports for any known class, it is very likely categorized as an unknown class. This behavior stems from Paper Eq.10 which ensures $\mathcal{A}^{y_{i}*}\neq\mathcal{A}^{y_{j}*}$ through constraining $\gamma^{y_{i}}\neq\gamma^{y_{j}}$. L-Reg further enhances the model’s ability to identify minimal supports for unknown classes by filtering out co-covariant features associated with other classes and thus generalizing to unseen domains. Therefore, the very interpretable features for unknown classes from unseen domains can be extracted using L-Reg. Fig. 5 (right side) demonstrates that the model with L-Reg can even extract facial features for the unknown person class and can generalize this to the unseen domain. Similarly, here we obtain (informal) atomic formulas as $h(\text{has a face},\text{is person},d\in D)\to\text{True}$, h(not $h(\text{not has a face},\text{is person},d\in D)\to\text{False}$.
@@ Appendix B Details of the logical framework for visual classification task
$\mathcal{L}=\left\langle F_{\mathcal{L}},M_{\mathcal{L}},\models_{\mathcal{L}},mng_{\mathcal{L}},\vdash_{\mathcal{L}}\right\rangle,$ (16)
@@
    Any bijection $f:P\textrightarrow Q$ that extends to a bijection between the tautological formula algebras of $\mathcal{L}^{P}$ and $\mathcal{L}^{Q}$ induces an isomorphism between $\mathcal{L}^{P}$ and $\mathcal{L}^{Q}$.
@@ Appendix E Apply L-Reg to ERM Baseline for mDG
To further validate L-Reg’s efficacy for mDG, we use ERM as the baseline on the TerraIncognita dataset. For a fair comparison, all experiments share the same hyperparameter settings and use the Regnety-16gf backbone. Original ERM results are also included alongside our reproduced results. The results in Table 8 reveal that ERM with L-Reg significantly improves mDG performance (from 49.9% to 52.9%).
@@ Appendix F Compare L-Reg with more regularization terms
We also compare L-Reg with other regularization terms: The Ortho-Reg - the orthogonality regularization that constrains the independence of each dimension of the semantic feature $z$; and Sparsity - implemented as Bernoulli Sample of the latent features from the sparse linear concept discovery models [41] on our used PIM backbone. To investigate this fairly, we re-implemented the Bernoulli Sample of the latent features from the Sparse Linear Concept Discovery Models [41] on the same PIM backbone that we used, to achieve the sparsity. Table 6&Table 7 demonstrate that L-Reg outperforms Ortho-Reg and Sparsity.
Especially, while a common sparse concept model may be able to achieve $\gamma^{y}\psi=z^{y}$ by filtering irrelevant features through the sparsity, it may not ensure $\gamma^{y_{i}}\neq\gamma^{y_{j}}$, which is crucial for disentangling features used for predicting different classes. This limitation can potentially lead to degradation in generalization performance for common sparse concept models. 6&7 indicate that while L-Reg consistently achieves overall improvement, the sparse concept-based approach does not consistently improve generalization, validating the aforementioned difference.
@@ Appendix G Limitation of L-Reg and possible solutions
As analyzed and discussed in the paper, L-Reg is based on the precondition that each dimension of the latent features represents an independent semantic.
We hypothesize this is due to the fact that our L-Reg is derived based on the precondition that $z^{i},z^{j}\in z,I\neq j$ is independent of each other. This condition holds for most deep-layer features but may not apply to shallow layers. Thus, applying L-Reg to the semantic features from the deep layers may improve the performance for unknown classes without negatively impacting known classes.
Derived from this hypothesis, another possible solution is further regularizing the independence, which may lead to further improvements. To validate this hypothesis, we test L-Reg by reinforcing independence with Ortho-Reg. MDG results in Table 8 and GCD results in Table 6&Table 7 show that combining L-Reg with Ortho-Reg leads to further improvements, whereas Ortho-Reg alone may not guarantee improvements. These findings support our hypothesis and suggest that L-Reg, particularly when applied to deep layers or in conjunction with Ortho-Reg, is beneficial. This suggests a direction for future work.
@@
Table 8: Results of mDG: Results of using ERM as the baseline. We use the ERM method as the baseline to test L-Reg’s efficacy. Ortho-Reg: orthogonality regularization. This table includes results: (1) The improved performance of L-Reg on ERM baseline. (2) Comparison between L-Reg and the Ortho-Reg on ERM baseline. (3) Using L-Reg and Ortho-Reg together yields further promotion, validating our ‘improper $z$’ hypothesis in the Paper limitation part. The used dataset is TerraIncognita. All experiments share the same hyperparameters except the added regularization term. Each group of experiments is run with seeds [0,1,2], and the averaged results for each domain and additionally with the standard deviation (Std) are reported.
@@ H.1 Multi-domain generalization
Datasets. We use PACS (4 domains, 9,991 samples, $7$ classes) [32], VLCS ($4$ domains, $10,729$ samples, $5$ classes) [18], OfficeHome (4 domains, 15,588 samples, 65 classes) [55], TerraIncognita (TerraIncognita, $4$ domains, $24,778$ samples, $10$ classes) [7], and DomainNet (6 domains, 586,575 samples, 345 classes) [42].
@@
TerraIncognita Location 100 Location 38 Location 43 Location 46 Avg.
@@ H.3 Combination of multi-domain generalization and generalized category discovery
PACS HomeOffice VLCS TerraIncognita DomainNet
@@
Table 23: MDG+GCD results: accuracy scores of each domain in TerraIncognita dataset.
TerraIncognita Avg art_painting cartoon photo sketch